OpenStudy (anonymous):

Suppose that a guy wire is located 15 ft up a telephone pole and is anchored to the ground 10 ft from the base of the pole. How long is the guy wire?

OpenStudy (anonymous):



OpenStudy (jlastino):

use the Pythagorean theorem: a^2+b^2=c^2 a=10 b=15 sub values and solve for c

OpenStudy (anonymous):

325? isnt that too long?

OpenStudy (jlastino):

remember that 325= c^2 take the square root 325 to find c :D